OpenStudy (anonymous):

Which expression is equivalent to cos(4x) + cos(2x)? 2cos(3x)cos x -2sin(3x)sin x 2sin(3x)sin x -2cos(3x)cos x

OpenStudy (goformit100):

Expand cos(4x) + cos(2x)

OpenStudy (ybarrap):

You can probably use these identities $$ \cos 4x=\cos (2\times2x) $$ But also $$ \cos 2x = \cos^2x - \sin^2x\\ \cos^2x=1-\sin^2x $$

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

is it 2cos(3x)cos x?

OpenStudy (anonymous):

this is ridiculous

OpenStudy (anonymous):

Refer to the attached Mathematica plots.

OpenStudy (triciaal):

one approach let 2A =B so you have cos 2B + cos B cos n theta = 2 cos theta (n-1) cos theta - cos (n-2) theta here n = 2 and theta = B what do you get ? next cos P + cos Q = 2 cos (P + Q )/2 * cos (P - Q)/2 substitute and simplify

OpenStudy (triciaal):